# Majorana Tensor Decomposition: A unifying framework for decompositions of fermionic Hamiltonians to Linear Combination of Unitaries

### v2

@@ I Introduction
Here, $\hat{U}_{k}$ are unitary operators, $u_{k}$ are complex coefficients, and $K$ is the total number of unitaries. By estimating the phase of the walker operator, the eigen-values of $\hat{H}$ can be extracted from $\textrm{arccos}(E_{n}/\lambda)$. The accuracy associated with the energy estimation is upper-bounded by $\epsilon\leq\lambda\epsilon_{W}$, where $\epsilon_{W}$ is the walker operator phase estimation accuracy. Consequently, employing an optimal phase estimation algorithm with Heisenberg scaling $\tilde{\mathcal{O}}(1/\epsilon_{W})$ [5, 9], the algorithm’s cost to achieve a target energy accuracy of $\epsilon$ would require $\tilde{\mathcal{O}}(\lambda/\epsilon)$ calls to the walker operator. It is evident that the cost of the LCU-based energy estimation method depends on both the 1-norm $\lambda$ and the implementation cost of the walker operator.
@@ II Majorana Tensor Decomposition Framework
where $\Omega_{\vec{w},\vec{m}}$ is a tensor with possibly complex coefficients, $\hat{p}_{\vec{w}}^{(\nu)}$ is a unitary operator expressed as a polynomial of Majorana operators of degree $\leq 4$ with a minimal number of terms, $\hat{V}^{(\nu)}_{\vec{m}}\equiv e^{\sum_{p>q}\theta^{(\vec{m},\nu)}_{pq}\hat{\gamma}_{p}\hat{\gamma}_{q}}\in\textrm{Spin}(4N)$ are unitary rotations that conserve the degree of $\hat{p}_{\vec{w}}^{(\nu)}$. For all decompositions shown in this work only spacial orbital rotations are used, which form a subgroup of $\textrm{Spin}(4N)$
@@
For example, in the linear case $\hat{p}_{1}^{\nu}=\gamma_{1\sigma,x}$, where $x=\{0,1\}$, all other linear in Majorana unitary operators can be obtained by conjugating with $\hat{U}_{\vec{\theta}}$ these two polynomials. For higher polynomial degrees, there is a still relatively low number of distinct $\hat{p}_{\vec{w}}^{(\nu)}$’s that give non-overlapping sets of unitary operators upon $\hat{U}_{\vec{\theta}}$ conjugation. We constrain the product $\prod_{\nu=1}^{L}\hat{p}_{\vec{w}}^{(\nu)}$ to yield a Majorana polynomial of degree $\leq 4$ for all $w_{1},...,w_{L}$ appearing in the sum, which restricts $L\leq 4$.
@@ II.1 MTD-$L^{4}$ decomposition
The idea behind the MTD-$L^{4}$ decomposition is representing unitaries as products of four unitaries that are first degree polynomials of Majoranas, thus having $L=4$. A way of writing this decomposition that conserves the spin-symmetry in Eq. (1) is
@@ II.1.3 SVD-based MTD-$L^{4}$
$=\sum_{\alpha_{1}=1}^{N}U_{i\alpha_{1}}^{(1)}S^{(1)}_{\alpha_{1}}\sum_{\alpha_{2}=1}^{N}U_{j\alpha_{2}}^{(2)}S^{(2,\alpha_{1})}_{\alpha_{2}}\times\cdots$
×$\ \ \cdots\times\sum_{\alpha_{3}=1}^{N}U_{k\alpha_{3}}^{(3)}S^{(3,\alpha_{1}\alpha_{2})}_{\alpha_{3}}V^{(3,\alpha_{1}\alpha_{2}\alpha_{3})}_{l}$ (33)
@@
$\hat{q}^{(1)}_{\alpha_{1}\alpha_{2}\alpha_{3},\sigma}$ $=\sum_{i}U_{i\alpha_{1}}^{(1)}\hat{\gamma}_{i\sigma,0}$ (34)
$\hat{q}^{(2)}_{\alpha_{1}\alpha_{2}\alpha_{3},\sigma}$ $=\sum_{j}U_{j\alpha_{2}}^{(2,\alpha_{1})}\hat{\gamma}_{j\sigma,1}$ (35)
@@ III.2 Fermionic LCU decompositions
Next, we illustrate how fermionic-based LCU decompositions can be expressed as MTD. Unitaries in these LCUs are usually constructed as orbital rotations acting on reflections $\hat{Q}_{ii\sigma}\equiv i\hat{\gamma}_{i\sigma,0}\hat{\gamma}_{i\sigma,1}=\hat{1}-2\hat{n}_{i\sigma}$, where $\hat{n}_{i\sigma}\equiv\hat{a}_{i\sigma}^{\dagger}\hat{a}_{i\sigma}$ represents the occupation number operator. The fermion-to-qubit mappings transform $Q_{ii\sigma}$ operator to the qubit operator $\hat{z}_{i\sigma}$.
@@ III.2.1 Single factorization
where $W_{ij}^{(\mathcal{l})}$ represents the Cholesky decomposition of the $g_{ijkl}$ tensor written as a $N^{2}\times N^{2}$ matrix $g_{ij;kl}$, and $c_{\rm{SF}}$ is a constant. We note that the $\mathcal{l}$ index is usually truncated at some $M\ll N^{2}$. Equation (51) illustrates how this decomposition can be cast as an MTD that separates $\hat{H}_{1}$ and $\hat{H}_{2}$, with $L=2$ [Eq. (10)]. The associated sum over $\vec{m}$ has trivial rotations $\hat{V}=\hat{1}$ and $M$ elements. The Majorana polynomials $\hat{p}$’s correspond to the Majorana products in Eq. (51). Additionally, the complete-square structure of the two-electron elements allows them to be expressed as squares of $\hat{p}_{\mathcal{l}}\equiv\sum_{\sigma}\sum_{ij}W_{ij}^{(\mathcal{l})}\hat{\gamma}_{i\sigma,0}\hat{\gamma}_{j\sigma,1}$ operators. Complete-squares can be encoded through the use of oblivious amplitude amplification, which implements through qubitization the second-degree Chebyshev polynomial of the normalized $\hat{p}_{\mathcal{l}}$, i.e. $2\left(\hat{p}_{\mathcal{l}}/N_{\mathcal{l}}^{(\rm{SF})}\right)^{2}-\hat{1}$, where the normalization constant $N_{\mathcal{l}}^{(\rm{SF})}\equiv 2\sum_{ij}|W_{ij}^{(\mathcal{l})}|$. Finally, we note that the original procedure did not diagonalize the one-electron term. The MTD representation of single factorization becomes
@@ III.2.3 Cartan sub-algebra
with $L=2$ [Eq. (10)]. The associated $\Omega$ tensor has an implicit Kronecker delta in the $m_{1}$ and $m_{2}$ indices. Despite being more flexible than DF, the non-linear optimization procedures required to obtain the Cartan sub-algebra decomposition make its classical computational cost expensive for large systems.
@@ III.2.4 Tensor Hypercontraction
$\ \ +\sum_{\mu\nu}\zeta_{\mu\nu}\sum_{\sigma}\hat{U}_{\mu}^{\dagger}\hat{Q}_{11\sigma}\hat{U}_{\mu}\sum_{\tau}\hat{U}_{\nu}^{\dagger}\hat{Q}_{11\tau}\hat{U}_{\nu},$ (58)
which corresponds to an $L=2$ MTD with $\vec{w}=1$ index [Eq. (10)]. Since the $\hat{U}$’s orbital rotations are only acting on orbital $1$, their implementation requires $\mathcal{O}(N)$ Givens rotations, as opposed to the $\mathcal{O}(N^{2})$ required for rotating all orbitals [24]. To better understand this reduction in the number of Givens rotations, we start by noting that for an arbitrary rotation $U\in\textrm{Spin}(N)$, the maximal torus theorem [56] allows us to write
@@
The corresponding MTD form of the THC decomposition is
$\hat{H}$ $=c_{\rm{THC}}\hat{1}+i\sum_{j}\mu_{j}\sum_{\sigma}\hat{U}_{0}^{\dagger}\hat{\gamma}_{j\sigma,0}\hat{\gamma}_{j\sigma,1}\hat{U}_{0}$
$\ \ +\sum_{\mu\nu}\zeta_{\mu\nu}\sum_{\sigma}\hat{U}_{\mu}^{\dagger}\hat{\gamma}_{1\sigma,0}\hat{\gamma}_{1\sigma,1}\hat{U}_{\mu}\sum_{\tau}\hat{U}_{\nu}^{\dagger}\hat{\gamma}_{1\tau,0}\hat{\gamma}_{1\tau,1}\hat{U}_{\nu}.$ (61)
@@ IV Discussion
The most time-consuming part of the fault-tolerant QPE algorithm is performing T-gates and their error correction. Considering that the number of applications of the Hamiltonian LCU oracle is proportional to the 1-norm of the used LCU, as a quantum resource cost we use a product of the LCU 1-norm by the number of T-gates required for an implementation of the Hamiltonian oracle. The latter corresponds to the number of T-gates in one $\mathtt{SELECT}$ and two $\mathtt{PREPARE}$ operations. We will refer to the quantum resource cost simply as hardness. Additionally, we note that optimizing the compilation strategy for the $\mathtt{SELECT}$ circuit is imperative for minimizing the hardness. Compilation of these circuits typically exploit structures present in the used LCU ansatz, as seen in Ref. 20 for DF and THC approaches which make use of the spin-symmetry that is present (i.e. working with $\hat{F}^{i}_{j}$ operators). Whether better circuits can be found for the decompositions that have been introduced in this work is an open question. We note that results for the THC are not shown in this work due to the computational difficulty of performing the associated tensor decomposition in a robust and efficient way.
@@
To assess efficiency of various LCU decompositions we present quantum resource estimates for small molecules (Table 2) and 1D hydrogen chains (Fig. 5 and Table 3), computational details are provided in Appendix B. Most of LCU methods can be partitioned into three categories based on the small molecule results: 1) low T-gate count and medium 1-norms, Pauli and Pauli-OO; 2) high T-gate count and low 1-norms, AC, OO-AC, and DF; 3) medium T-gate count and high 1-norms, $L^{4}$-CP4 and $L^{4}$-MPS. The only method that combined low 1-norms and low T-gate counts is $L^{4}$-SVD.
Molecule $\Delta E/2$ Pauli OO-Pauli AC OO-AC DF $L^{4}$-SVD $L^{4}$-CP4 $L^{4}$-MPS
H2 1.68 2.54(632) 2.54(632) 2.18(1040) 2.18(1040) 2.11(3792) 2.54(3264) 19.2(3342) 3.80(4658)
LiH 7.72 13.7(1952) 13.0(1944) 10.2(95895) 10.2(93570) 9.78(23671) 10.9(12086) 252.5(9944) 67.9(12226)
BeH2 16.0 26.4(2144) 25.7(2144) 20.6(99055) 20.6(98214) 20.0(31529) 22.7(15488) 296.3(11796) 144.9(14284)
H2O 61.5 89.2(2504) 77.4(2480) 71.2(166249) 70.0(166249) 68.9(31529) 72.7(15488) 367.0(11796) 279.5(14284)
NH3 51.9 84.8(5016) 70.6(3728) 63.0(585623) 61.4(353231) 60.3(39550) 64.5(19548) 889.3(14446) 344.1(16664)
@@
Table 2: 1-norms for electronic structure Hamiltonians in the STO-3G basis [57], the numbers of T-gates required to implement the associated Hamiltonian oracle circuit are given in parenthesis. $\Delta E/2$ corresponds to a lower bound for the 1-norm. Pauli: Pauli products as unitaries; AC: anti-commuting grouping technique; OO-prefix indicates that orbitals were optimized to minimize the 1-norm; DF: double factorization; $L^{4}$-SVD, $L^{4}$-CP4, and $L^{4}$-MPS: MTD-$L^{4}$ decompositions introduced in Sec. II.1.1.
@@
Regarding the scaling of hardness with the system size (Table 3), the only decomposition which presented a better scaling than Pauli-OO is $L^{4}$-CP4 LCU. We note that extremely large 1-norms were obtained for the $L^{4}$-CP4 tensor decomposition for the small molecules (Table 2). We attribute this to an implementation that was used for the CP4 decomposition, and having a more robust implementation will be necessary for larger systems, as discussed in Ref. 58. At the same time, we can see how the $L^{4}$-SVD decomposition gave competitive 1-norms for these molecules, having that both the $L^{4}$-SVD and the $L^{4}$-CP4 decompositions correspond to the same $L^{4}$ LCU. On the other hand, the SVD-based 1-norms for the hydrogen chains were larger than their CP4 counterparts, having that the SVD decomposition became prohibitively expensive to compute for the larger chains. Overall, these results establish the $L^{4}$ LCU decomposition as a very promising decomposition in terms of hardness; finding the corresponding fragments in a robust and efficient way is work in progress. We believe that this decomposition could be useful for hybrid LCU approaches by obtaining a few large MTD-$L^{4}$ fragments before decomposing the rest of the Hamiltonian with some other method. However, exploring these kind of hybrid LCUs is outside the scope of this work.
@@
The anti-commuting (AC) decomposition, when using the orbital optimized Hamiltonians, yielded very competitive 1-norms and requires the fewest qubits. However, the large number of rotations required for its $\mathtt{SELECT}$ circuit make it require an extremely large number of T-gates, making its hardness one of the worst out of all the studied LCUs. As such, we believe that an implementation of the AC decomposition which loads the Givens rotation angles in the quantum computer could lower significantly the cost compared to the current implementation that we proposed where angles are “hard coded” into the circuits.
@@
The results for the DF shown in this work correspond to the implementation discussed in the Appendix A. The “optimized DF” results in Fig. 5 and Table 3 correspond to the DF implementation discussed in Ref. 20 where each Toffoli gate has been considered to require 4 T-gates [59]. We note how this decomposition also presents one of the best hardness scalings. However, the number of qubits required for the optimized DF is an order of magnitude larger than that required by other approaches.
The classical pre-processing cost: The Pauli products decomposition of a Hamiltonian has $\mathcal{O}(N^{4})$ operators. However, orbital localization schemes can significantly reduce this to $\mathcal{O}(N^{2})$ for large molecules [60]; the cost for optimized orbital localization schemes typically scales as $\mathcal{O}(N^{3})$ [31]. The sorted insertion algorithm, used in AC, has a sorting cost of $\mathcal{O}(M^{2})$, where $M$ is the number of elements in the list being sorted. This results in a decomposition cost of $\mathcal{O}(N^{8})$ for AC and $\mathcal{O}(N^{4})$ when grouping is done on localized orbitals. For the fermion-based methods, the cost for DF is dominated by the diagonalization of the reshaped $N^{2}\times N^{2}$ tensor $g_{ij;kl}$, which scales as $\mathcal{O}(N^{6})$. For the $L^{4}$-MPS, the cost for each step in the MPS factorization (see Fig. 3) goes as $\mathcal{O}(N^{5})$, thus being the overall scaling cost of this decomposition. The $L^{4}$-SVD decomposition quickly becomes prohibitively expensive to do as the system size grows, while the $L^{4}$-CP4 decomposition requires $\mathcal{O}(N^{3}W)$, where $W$ is the number of terms in the decomposition [58]. Among the algorithms presented in this work, orbital-localized AC has the best theoretical scaling, followed by $L^{4}$-MPS. However, in practice, the fermionic-based decompositions (DF and $L^{4}$-MPS) tend to have lower prefactors due to their efficient utilization of linear algebra routines. For example, in the case of the linear H30 chain, orbital-localized AC took approximately 5 minutes, while DF required around 0.1 seconds and $L^{4}$–MPS required around 0.3 seconds. These calculations were done using a single core of a 2.2 GHz 6-Core Intel i7 processor.
In summary, this work has introduced the MTD framework, which provides a unified approach to LCU decompositions of the electronic structure Hamiltonian. This framework not only presents a unifying perspective but also leads to the MTD-$L^{4}$ decomposition, which expresses the electronic structure tensor as a minimal-rank representation. This decomposition establishes a connection between popular tensor decompositions and LCUs. A study over small systems revealed the LCU based on Pauli products with an orbital optimization to be the most competitive. However, the MTD-$L^{4}$ decomposition had better scaling coefficients, suggesting that as we go to larger systems this decomposition could be more efficient to implement.

### v3

@@ I Introduction
Here, $\hat{U}_{k}$ are unitary operators, $u_{k}$ are complex coefficients, and $K$ is the total number of unitaries. By estimating the phase of the walk operator, the eigenvalues of $\hat{H}$ can be extracted from $\textrm{arccos}(E_{n}/\lambda)$. The accuracy associated with the energy estimation is upper-bounded by $\epsilon\leq\lambda\epsilon_{W}$, where $\epsilon_{W}$ is the walk operator phase estimation accuracy. Consequently, employing an optimal phase estimation algorithm with Heisenberg scaling $\tilde{\mathcal{O}}(1/\epsilon_{W})$ [5, 9], the algorithm’s cost to achieve a target energy accuracy of $\epsilon$ would require $\tilde{\mathcal{O}}(\lambda/\epsilon)$ calls to the walk operator. It is evident that the cost of the LCU-based energy estimation method depends on both the 1-norm $\lambda$ and the implementation cost of the walk operator.
@@ II Majorana Tensor Decomposition Framework
where $\Omega_{\vec{w},\vec{m}}$ is a tensor with possibly complex coefficients, $\hat{p}_{\vec{w}}^{(\nu)}$ is a unitary operator expressed as a polynomial of Majorana operators of degree $\leq 4$ with a minimal number of terms, $\hat{V}^{(\nu)}_{\vec{m}}\equiv e^{\sum_{p>q}\theta^{(\vec{m},\nu)}_{pq}\hat{\gamma}_{p}\hat{\gamma}_{q}}\in\textrm{Spin}(4N)$ are unitary rotations that conserve the degree of $\hat{p}_{\vec{w}}^{(\nu)}$. We also constrain the product $\prod_{\nu=1}^{L}\hat{p}_{\vec{w}}^{(\nu)}$ to yield a Majorana polynomial of degree $\leq 4$ for all $w_{1},...,w_{L}$ appearing in the sum, which restricts $L\leq 4$. For all decompositions shown in this work only spacial orbital rotations are used, which form a subgroup of $\textrm{Spin}(4N)$
@@
For example, in the linear case $\hat{p}_{1}^{\nu}=\gamma_{1\sigma,x}$, where $x=\{0,1\}$, all other linear in Majorana unitary operators can be obtained by conjugating with $\hat{U}_{\vec{\theta}}$ these two polynomials. For higher polynomial degrees, there is a still relatively low number of distinct $\hat{p}_{\vec{w}}^{(\nu)}$’s that give non-overlapping sets of unitary operators upon $\hat{U}_{\vec{\theta}}$ conjugation.
@@ II.1 MTD-$L^{4}$ decomposition
Here we start with a novel LCU decomposition, MTD-$L^{4}$, that has not been suggested before. The idea behind the MTD-$L^{4}$ decomposition is representing unitaries as products of four unitaries that are first degree polynomials of Majoranas, thus having $L=4$. A way of writing this decomposition that conserves the spin-symmetry in Eq. (1) is
@@ II.1.3 SVD-based MTD-$L^{4}$
$=\sum_{\alpha_{1}=1}^{N}U_{i\alpha_{1}}^{(1)}S^{(1)}_{\alpha_{1}}\sum_{\alpha_{2}=1}^{N}U_{j\alpha_{2}}^{(2)}S^{(2,\alpha_{1})}_{\alpha_{2}}$
×$\ \ \times\sum_{\alpha_{3}=1}^{N}U_{k\alpha_{3}}^{(3)}S^{(3,\alpha_{1}\alpha_{2})}_{\alpha_{3}}V^{(3,\alpha_{1}\alpha_{2}\alpha_{3})}_{l}$ (33)
@@
$\hat{q}^{(1)}_{\alpha_{1},\sigma}$ $=\sum_{i}U_{i\alpha_{1}}^{(1)}\hat{\gamma}_{i\sigma,0}$ (34)
$\hat{q}^{(2)}_{\alpha_{1}\alpha_{2},\sigma}$ $=\sum_{j}U_{j\alpha_{2}}^{(2,\alpha_{1})}\hat{\gamma}_{j\sigma,1}$ (35)
@@ III.2 Fermionic LCU decompositions
Next, we illustrate how fermionic-based LCU decompositions can be expressed as MTD. Unitaries in these LCUs are usually constructed as orbital rotations acting on reflections $\hat{Q}_{ii\sigma}\equiv i\hat{\gamma}_{i\sigma,0}\hat{\gamma}_{i\sigma,1}=2\hat{n}_{i\sigma}-\hat{1}$, where $\hat{n}_{i\sigma}\equiv\hat{a}_{i\sigma}^{\dagger}\hat{a}_{i\sigma}$ represents the occupation number operator. The fermion-to-qubit mappings transform $Q_{ii\sigma}$ operator to the qubit operator $\hat{z}_{i\sigma}$.
@@ III.2.1 Single factorization
where $W_{ij}^{(\mathcal{l})}$ represents the Cholesky decomposition of the $g_{ijkl}$ tensor written as a $N^{2}\times N^{2}$ matrix $g_{ijkl}$, and $c_{\rm{SF}}$ is a constant. We note that the $\mathcal{l}$ index is usually truncated at some $M\ll N^{2}$. Equation (51) illustrates how this decomposition can be cast as an MTD that separates $\hat{H}_{1}$ and $\hat{H}_{2}$, with $L=2$ [Eq. (10)]. The associated sum over $\vec{m}$ has trivial rotations $\hat{V}=\hat{1}$ and $M$ elements. The Majorana polynomials $\hat{p}$’s correspond to the Majorana products in Eq. (51). Additionally, the complete-square structure of the two-electron elements allows them to be expressed as squares of $\hat{p}_{\mathcal{l}}\equiv\sum_{\sigma}\sum_{ij}W_{ij}^{(\mathcal{l})}\hat{\gamma}_{i\sigma,0}\hat{\gamma}_{j\sigma,1}$ operators. Complete-squares can be encoded through the use of oblivious amplitude amplification, which implements through qubitization the second-degree Chebyshev polynomial of the normalized $\hat{p}_{\mathcal{l}}$, i.e. $2\left(\hat{p}_{\mathcal{l}}/N_{\mathcal{l}}^{(\rm{SF})}\right)^{2}-\hat{1}$, where the normalization constant $N_{\mathcal{l}}^{(\rm{SF})}\equiv 2\sum_{ij}|W_{ij}^{(\mathcal{l})}|$. Finally, we note that the original procedure did not diagonalize the one-electron term. The MTD representation of single factorization becomes
@@ III.2.3 Cartan sub-algebra
with $L=2$ [Eq. (10)]. The associated $\Omega$ tensor has an implicit Kronecker delta in the $m_{1}$ and $m_{2}$ indices. Despite being more flexible than DF, the non-linear optimization procedure required to obtain the Cartan sub-algebra decomposition becomes computationally expensive for large systems.
@@ III.2.4 Tensor Hypercontraction
$\ \ +\sum_{\mu,\nu=1}^{K}\zeta_{\mu\nu}\sum_{\sigma}\hat{U}_{\mu}^{\dagger}\hat{Q}_{11\sigma}\hat{U}_{\mu}\sum_{\tau}\hat{U}_{\nu}^{\dagger}\hat{Q}_{11\tau}\hat{U}_{\nu},$ (58)
which corresponds to an $L=2$ MTD with $\vec{w}=1$ index [Eq. (10)], the limit $K$ is referred as the rank of the THC decomposition. Since the $\hat{U}$’s orbital rotations are only acting on orbital $1$, their implementation requires $\mathcal{O}(N)$ Givens rotations, as opposed to the $\mathcal{O}(N^{2})$ required for rotating all orbitals [24]. To better understand this reduction in the number of Givens rotations, we start by noting that for an arbitrary rotation $U\in\textrm{Spin}(N)$, the maximal torus theorem [56] allows us to write
@@ IV Results and Discussion
The most time-consuming part of the fault-tolerant QPE algorithm is performing T-gates and their error correction. Considering that the number of applications of the Hamiltonian LCU oracle is proportional to the 1-norm of the used LCU, as a quantum resource cost we use a product of the LCU 1-norm by the number of T-gates required for an implementation of the Hamiltonian oracle. The latter corresponds to the number of T-gates in one $\mathtt{SELECT}$ and two $\mathtt{PREPARE}$ operations. We will refer to the quantum resource cost simply as hardness. Additionally, we note that optimizing the compilation strategy for the $\mathtt{SELECT}$ circuit is imperative for minimizing the hardness. Compilation of these circuits typically exploit structures present in the used LCU ansatz, as seen in Ref. 20 for DF and THC approaches which make use of the spin-symmetry that is present (i.e. working with $\hat{F}^{i}_{j}$ operators). Whether better circuits can be found for the decompositions that have been introduced in this work is an open question.
All our fermionic LCU decompositions are performed on the two-body part of the Hamiltonian, which we define as the part of the Hamiltonian quartic in Majorana operators. Since we know that the remaining one-body part of the Hamiltonian can be optimally treated using DF [23], all 1-norms corresponding to fermionic LCU decompositions reported in this paper are, therefore, sum of the spectral range of the one-body part and the 1-norm for the fermionic LCU decomposition of the two-body part. We have obtained the THC decompositions using a non-linear optimization performed with the Levenberg-Marquardt algorithm [57, 58], details of our THC implementation are in Appendix B. The Toffoli gate counts for our THC decompositions are obtained using Openfermion [59] and the reported T-gate counts are 4 times the Toffoli counts [60].
@@
To assess efficiency of various LCU decompositions we present quantum resource estimates for small molecules (Table 2) and 1D hydrogen chains (Fig. 5 and Table 3), computational details are in Appendix B. Most of LCU methods can be partitioned into four categories based on the small molecule results: 1) low T-gate count and medium 1-norms, Pauli and Pauli-OO; 2) medium 1-norms and medium T-gate count, $L^{4}$-SVD and THC; 3) high T-gate count and low 1-norms, AC, OO-AC, and DF; 4) medium T-gate count and high 1-norms, $L^{4}$-CP4 and $L^{4}$-MPS.
Molecule H2 LiH BeH2 H2O NH3
$\Delta E/2$ 0.815 4.93 9.99 41.9 33.8
Pauli 1.58(632) 13.0(1952) 22.8(2144) 71.9(2512) 69.2(5032)
OO-Pauli 1.58(632) 12.4(1936) 21.9(2136) 61.0(2488) 54.5(3744)
AC 1.49(1040) 10.2(95362) 18.0(98522) 57.2(166700) 48.9(588301)
OO-AC 1.49(1040) 10.2(92715) 17.9(96860) 55.7(162665) 46.8(355022)
DF 1.37(3792) 9.34(23671) 16.4(31529) 53.7(31529) 44.7(39550)
THC 1.8 (2500) 11.18 (5976) 19.59 (6832) 58.55 (6832) 49.26 (8316)
$L^{4}$-SVD 2.19 (3264) 11.8 (12086) 22.0 (15488) 61.1 (15488) 53.3 (19548)
$L^{4}$-CP4 5.65(7556) 34.5 (13044) 61.1 (14416) 125.0 (14416) 121.0 (15788)
$L^{4}$-MPS 4.63(3750) 130.08(12384) 279.27(14432) 474.1(14284) 612.45(16664)
@@
Table 2: 1-norms for electronic structure Hamiltonians in the STO-3G basis [61], the numbers of T-gates required to implement the associated Hamiltonian oracle circuit are given in parenthesis. $\Delta E/2$ corresponds to a lower bound for the 1-norm. Pauli: Pauli products as unitaries; AC: anti-commuting grouping technique; OO-prefix indicates that orbitals were optimized to minimize the 1-norm; DF: double factorization; THC: tensor hypercontraction; $L^{4}$-SVD, $L^{4}$-CP4, and $L^{4}$-MPS: MTD-$L^{4}$ decompositions introduced in Sec. II.1.1.
@@
Regarding the scaling of hardness with the system size (Table 3), the only decompositions which presented a better scaling than Pauli-OO are $L^{4}$-CP4 and THC. The THC decomposition offers the most optimal scaling with increasing system size $N$, hardness scaling of $O(N^{2.19})$ is obtained empirically from results of Table 3. Thus, THC can be advantageous over Pauli-OO for systems with $N>23$. We note that extremely large 1-norms were obtained for the $L^{4}$-CP4 tensor decomposition for the small molecules (Table 2). We attribute this to an implementation that was used for the CP4 decomposition, and having a more robust implementation will be necessary for larger systems, as discussed in Ref. 62. At the same time, $L^{4}$-SVD decomposition gave competitive 1-norms for these molecules. On the other hand, the SVD-based 1-norms for the hydrogen chains were larger than their CP4 counterparts, which makes the SVD decomposition prohibitively expensive for larger chains. Overall, these results establish the $L^{4}$ LCU decomposition as a promising decomposition in terms of hardness; finding the corresponding fragments in a robust and efficient way is work in progress. We believe that this decomposition could be useful for hybrid LCU approaches by obtaining a few large MTD-$L^{4}$ fragments before decomposing the rest of the Hamiltonian with some other method. However, exploring these kind of hybrid LCUs is beyond the scope of this work.
@@
The anti-commuting (AC) decomposition, when using the orbital optimized Hamiltonians, yielded very low 1-norms and requires the fewest qubits. However, the large number of rotations required for its $\mathtt{SELECT}$ circuit increases the T-gate count and makes its hardness one of the worst out of all the studied LCUs. As such, we believe that an implementation of the AC decomposition which loads the Givens rotation angles in the quantum computer could lower significantly the cost compared to the current implementation that we proposed where angles are “hard coded” into the circuits.
@@
The results for the DF method correspond to the implementation discussed in the Appendix A. The “optimized DF” results in Fig. 5 and Table 3 correspond to the DF implementation discussed in Ref. 20 where each Toffoli gate has been considered to require 4 T-gates [60]. We note that this decomposition also presents one of the best hardness scalings. However, the number of qubits required for the optimized DF is an order of magnitude larger than that required by most other approaches. Only the THC decomposition involves similar numbers of ancilla qubits as the Optimized DF approach.
@@
The classical pre-processing cost: The Pauli products decomposition of a Hamiltonian has $\mathcal{O}(N^{4})$ operators. However, orbital localization schemes can reduce the scaling to $\mathcal{O}(N^{2})$ for large molecules [63]; the cost for optimized orbital localization schemes typically scales as $\mathcal{O}(N^{3})$ [31]. The sorted insertion algorithm, used in AC, has a sorting cost of $\mathcal{O}(M^{2})$, where $M$ is the number of elements in the list being sorted. This results in a decomposition cost of $\mathcal{O}(N^{8})$ for AC and $\mathcal{O}(N^{4})$ when grouping is done on localized orbitals. For the fermionic methods, the cost for DF is dominated by the diagonalization of the reshaped $N^{2}\times N^{2}$ tensor $g_{ijkl}$, which scales as $\mathcal{O}(N^{6})$.
For our THC implementation, the computation of an approximate Hessian matrix for the update step in the Levenberg-Marquardt algorithm dominates the computational cost. Assuming that the rank of the THC decomposition, $K$ [Eq. (58)] scales linearly with system size, the computational cost of our THC implementation scales as $O(N^{8})$. This is in contrast to the $O(N^{6})$ scaling of Openfermion’s L-BFGS-B THC implementation. Our implementation of THC offers greater robustness with the choice of initial parameters in exchange for the higher computational scaling. For $L^{4}$-MPS, the cost for each step in the MPS factorization (see Fig. 3) goes as $\mathcal{O}(N^{5})$, thus being the overall scaling cost of this decomposition. The $L^{4}$-SVD decomposition quickly becomes prohibitively expensive to do as the system size grows, while the $L^{4}$-CP4 decomposition requires $\mathcal{O}(N^{3}W)$, where $W$ is the number of terms in the decomposition [62]. Among the algorithms presented in this work, orbital-localized AC has the best theoretical scaling, followed by $L^{4}$-MPS. However, in practice, the fermionic-based decompositions (DF and $L^{4}$-MPS) tend to have lower prefactors due to their efficient utilization of linear algebra routines. For example, in the case of the linear H30 chain, orbital-localized AC took approximately 5 minutes, while DF required around 0.1 seconds and $L^{4}$–MPS required around 0.3 seconds. These calculations were done using a single core of a 2.2 GHz 6-Core Intel i7 processor.
V Conclusion
In summary, this work has introduced the MTD framework, which provides a unified approach to LCU decompositions of the electronic structure Hamiltonian. This framework not only presents a unifying perspective but also leads to new decompositions like MTD $L^{4}$, the latter expresses the electronic structure tensor as a minimal-rank representation. MTD establishes a connection between popular tensor decompositions and LCUs. A study over small systems revealed the LCU based on Pauli products with an orbital optimization to be the most competitive. However, the MTD $L^{4}$-CP4 and THC decompositions had better scalings with the system size, suggesting that for larger systems these decompositions could be more efficient to implement.